The solid waste generated each year in the cities of the US is increasing. The solid waste generated, in millions of tons, was 82.3 in 1960 and 139.1 in 1980. The trend appears linear during this time.
A) Construct a formula for the amount of municipal solid waste generated in the US by finding the equation of the line through these two points.
B) Use this formula to predict the amount of municipal solid waste generated in the US in millions of tons in the year 2020.

Answer by tutorcecilia(2152)

m=y1-y2)/x1-x2) [use the slope formula]
(1960, 82.3) and (1980, 139.10)
Let 1960 = 1
Let 1980=22[identify two coordinate points]
(1, 82.3) and (22, 139.10)
m=(139.10-82.3)/22-1)=56.8/21=2.705[plug-in the values and solve for the slope]
.
A) Construct a formula for the amount of municipal solid waste generated in the US by finding the equation of the line through these two points.
y=mx+b [use the slope-intercept form of a line]
139.10=(2.705)(22)+b [plug-in the values for one point and the slope. Solve.]
139.10=59.51x+b
79.59=b
y=2.705x+79.59 [equation of the line]
.

B) Use this formula to predict the amount of municipal solid waste generated in the US in millions of tons in the year 2020.
Let 2020=63
y=2.705(63)+79.59 [plug-in the value of x=63 (year 2020) and solve for y]
y=250.01 million tons of waste will be generated in the year 2020!